# Tailoring anisotropy in kirigami metamaterial skins with pop-up folding hinges

### v1

@@ 4 Kinematic analysis and parametric exploration
In this section, we further explore the design space of kirigami with folding hinges that feature the desired global mechanism, to identify geometries that yield asymmetric pop-ups. This parametric study is guided by nonlinear kinematic formulas, that characterize the finite-amplitude mechanism motion of the selected unit cells. We begin by introducing our kinematic formulas for a generic unit cell and continue by performing the parametric study and identifying representative patterns to be further analyzed.
@@ 4.1 Kinematic analysis
Fig. 5: Example of kinematic simulation. (A-C) Three instances of the kinematic evolution of a asymmetric unit cell with mechanism, obtained by setting slit openings $\bar{w}_{x}=0.2$, $\bar{w}_{y}=0.1$, fold lengths ¯$\bar{\ell}_{1}=\bar{\ell}_{2}=\bar{\ell}_{3}=\bar{\ell}_{4}=0.3$ and fold angles $\lambda_{1}=\pi/4$, $\lambda_{2}=\pi/8$, $\lambda_{3}=\pi/12$, $\lambda_{4}=\pi/16$. The panel rotation angle $\theta$ is the only DOF; specifically $\theta=0$ in (A), $\theta=\pi/16$ in (B) and $\theta=\pi/32$ in (C). (D) Evolution of strains, normalized heights of the pop-ups and normalized opening areas of the pop-ups as a function of $\theta$, for the unit cell in (A).
To explain how we derive our kinematic formulas, we consider the pattern shown in Fig. 5A; its design parameters are reported in the figure’s caption. In particular, we choose all hinge lengths to be identical to preserve the global mechanism. To be as generic as possible, we select different values of slit opening $\bar{w}_{x}$ and $\bar{w}_{y}$, and we choose distinct values for all hinge angles $\lambda_{i}$. Two instances of the kinematic evolution of the selected unit cell as a function of the degree of freedom (DOF) $\theta$ are shown in Figs. 5B,C. Our kinematic formulas are obtained from a single panel of the unit cell; we choose the top-right one.
@@
As the panels in the unit counter-rotate by angle $\theta$, the above mentioned diagonals in the top-right panel rotate so that the angle between the diagonal of length $d_{x}$ and $x$ becomes $\alpha_{x}-\theta$, while the angle between $d_{y}$ and $y$ becomes $\alpha_{y}+\theta$. The projection of the diagonal of length $d_{x}$ onto the $x$ axis and that of $d_{y}$ onto $y$ can then be used to find the overall strains of the cell, as follows:

### v2

@@ 3 Kinematic analysis and parametric exploration
We now try to navigate the vast design space introduced in the previous section, with the goal of understanding which combinations of parameters preserve the global mechanism featuring panel counter-rotations, and identifying geometries that yield asymmetric pop-ups. We are particularly interested in understanding how the pop-up parameters relate to in-plane kinematics — an aspect that is unique to kirigami metamaterials with folding hinges.
As previously mentioned, the metamaterials we investigate here are variations of the rotating squares pattern. In  A, by analyzing the mechanisms and states of self stress of such patterns, we show that the presence of folding hinges does not affect the global mechanism. We also show that, in the case of asymmetric geometries, we cannot choose hinge lengths independently (we need ¯$\bar{\ell}_{1}=\bar{\ell}_{3}$ and ¯$\bar{\ell}_{2}=\bar{\ell}_{4}$), while hinge angles can be chosen independently. These considerations considerably restrict the available design space; however, the possibility of choosing hinge angles independently can yield interesting asymmetric designs, as shown below. Our conclusions echo recent findings on the need of having parallelogram-shaped slits to preserve the global mechanism of rotating squares metamaterials and their variations [74, 75, 73], including a formal theorem on this matter by Dang et al. that extends to patterns with even less symmetry than ours [76].
@@ 3.1 Kinematic analysis
We now probe this restricted design space to identify geometries that yield asymmetric pop-ups. This parametric study is guided by nonlinear kinematic formulas, that characterize the finite-amplitude mechanism motion of the selected unit cells. We begin by introducing our kinematic formulas for a generic unit cell and continue by performing the parametric study and identifying representative patterns to be further analyzed.
Fig. 3: Example of kinematic simulation. (a)-(c) Three instances of the kinematic evolution of a asymmetric unit cell with mechanism, obtained by setting slit openings $\bar{w}_{x}=0.2$, $\bar{w}_{y}=0.1$, fold lengths ¯$\bar{\ell}_{1}=\bar{\ell}_{2}=\bar{\ell}_{3}=\bar{\ell}_{4}=0.3$ and fold angles $\lambda_{1}=\pi/4$, $\lambda_{2}=\pi/8$, $\lambda_{3}=\pi/12$, $\lambda_{4}=\pi/16$. The panel rotation angle $\theta$ is the only DOF; specifically $\theta=0$ in (a), $\theta=\pi/16$ in (b) and $\theta=\pi/32$ in (c). (d) Evolution of strains, normalized heights of the pop-ups and normalized opening areas of the pop-ups as a function of $\theta$, for the unit cell in (a).
To explain how we derive our kinematic formulas, we consider the pattern shown in Fig. 3(a). In particular, we choose all hinge lengths to be identical to preserve the global mechanism. To be as generic as possible, we select different values of slit opening $\bar{w}_{x}$ and $\bar{w}_{y}$, and we choose distinct values for all hinge angles $\lambda_{i}$. Two instances of the kinematic evolution of the selected unit cell as a function of the degree of freedom (DOF) $\theta$ are shown in Figs. 3(b), (c). Our kinematic formulas are obtained by tracking the motion of a single panel, and we choose the top-right one.
@@
As the panels in the unit counter-rotate by angle $\theta$, the above mentioned diagonals in the top-right panel rotate so that the angle between the diagonal of length $d_{x}$ and $x$ becomes $\alpha_{x}-\theta$, while the angle between $d_{y}$ and $y$ becomes $\alpha_{y}+\theta$. The projection of the diagonal of length $d_{x}$ onto the $x$ axis and that of $d_{y}$ onto $y$ can then be used to find the overall strains of the cell,